what will be the Vaan der Waals equation for 1 mole of gas if the force of attraction between gas molecules is
negligible

Grade:11

2 Answers

Arun
6 years ago
Dear student
 
Vander walls equatiin is:
(P + n²*a/v²) (V-nb) = R*T
Now n = 1 mole
a = force of attraction = 0
Hence equatiin will be:
P(V-b) = RT
 
Regards
Arun (askIITians forum expert)
Aman
6 years ago
in Vander waal equation a stands for molecular force of attraction(P+an2/v2)( V-nb) =nRT n=1 and a=0 soP(V-b) =RT